Can someone help me with this?
You roll two fair dice, one red and one green.
a) what is the probability of getting a number less than 5 on both?
b) what is the probability of getting a sum of 9 on the two dice?
c) what is the probability of getting a 5 on both?
thank you
JIll
: Can someone help me with this?
You roll two fair dice, one red and one green.
a) what is the probability of getting a number less than 5 on both?
b) what is the probability of getting a sum of 9 on the two dice?
c) what is the probability of getting a 5 on both?
thank you
JIll

Answer by Fombitz(1755)

a.Less than 5: 1,2,3,4
P(<5)=4/6
Since each die is independent, the probabilities multiply,
P(<5 and <5)=4/6*4/6=16/36=4/9
b)Let's look at all of the possible die outcomes and the sum
1, 1, 2
1, 2, 3
1, 3, 4
1, 4, 5
1, 5, 6
1, 6, 7
2, 1, 3
2, 2, 4
2, 3, 5
2, 4, 6
2, 5, 7
2, 6, 8
3, 1, 4
3, 2, 5
3, 3, 6
3, 4, 7
3, 5, 8
3, 6, 9
4, 1, 5
4, 2, 6
4, 3, 7
4, 4, 8
4, 5, 9
4, 6, 10
5, 1, 6
5, 2, 7
5, 3, 8
5, 4, 9
5, 5, 10
5, 6, 11
6, 1, 7
6, 2, 8
6, 3, 9
6, 4, 10
6, 5, 11
6, 6, 12
And then the frequency of the sums
2,1
3,2
4,3
5,4
6,5
7,6
8,5
9,4
10,3
11,2
12,1
36 total die outcomes.
The sum 9 comes up 4 times.
P(sum 9 on 2 die roll)=4/36=1/9
c) P(5 on one) : 1/6
Both die are independent, multiply the probability.
P(5 on both)=1/6*1/6=1/36